What is the pH of the resulting solution when 22.43 mL 0.1472 M Ba(OH)2 is added to 48.39 mL 0.1623 M HCl? Enter your answer to three significant figures.
1
Expert's answer
2020-08-23T07:44:28-0400

moles of Ba(OH)2 = 22.43 × 0.1472 / 1000

= 3.30 × 10-3 mol.

moles of HCl = 48.39 × 0.1623 / 1000 .

= 7.85 × 10-3 mol.


1 mol .of Ba(OH)2 react with 2 mol of HCl .


So, Reacting mol of HCl = 6.60 × 10-3 .


Rest mol of HCl = 7.85 × 10-3 - 6.60 × 10-3


Rest mol of HCl = 1.25 × 10-3 .


Concentration of HCl =

1.25 × 10-3/ 77.82 × 10-3.


= 1.25 / 77.82 = 0.01606 M.


PH = - log concentration of H+.


PH = - log 0.01606


= - (-1.78 ) .


= + 1.78 .